# Separable commutative algebras in equivariant homotopy theory

### v1

@@ 7. Separably closed commutative algebras
Next, we record a few consequences of this condition.
@@
Let $\mathscr{F}$ be a family of subgroups of a finite group $G$, i.e. a possibly empty set of subgroups of $G$, closed under sub-conjugation. We denote by $\operatorname{Fin}_{G}/\mathscr{F}$ the full subcategory of $\operatorname{Fin}_{G}$ spanned by those finite $G$-sets $X$ which admits a (possibly empty) decomposition into orbits $X\simeq\coprod_{i}G/H_{i}$ with $H_{i}\not\in\mathscr{F}$. We also denote by $\operatorname{Fin}_{G}^{\mathrm{free}}$ the category of finite free $G$-sets. To simplify notation we we will write
@@
Let $G$ be a finite group and assume $A\in\operatorname{Fun}(BG,\operatorname{CAlg})$ satisfies IC and RC. Then, there is a unique factorization
${\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Perf}(A)^{hG})}$${\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Perf}(A)^{tG})}$${\mathcal{F}\mathrm{in}^{\operatorname{op}}_{G}\simeq\prod_{(H)}(\mathcal{F}\mathrm{in}_{W_{G}(H)}^{\mathrm{free}})^{\operatorname{op}}}$${\prod_{e\not=(H)}(\mathcal{F}\mathrm{in}_{W_{G}(H)}^{\mathrm{free}})^{\operatorname{op}},}$$\scriptstyle{q}$$\scriptstyle{A^{(-)}}$$\scriptstyle{\mathrm{proj}}$$\scriptstyle{\mathrm{\iota}}$
@@
Since the product in the bottom left corner is a product in the preadditive $\infty$-category of $\mathbb{E}_{\infty}$-monoids in spaces, it is also a coproduct (see [GGN15, Proposition 2.3]), and the indicated $\iota$ exists if and only if $q\circ A^{(-)}$ is zero on the factor corresponding to $(H)=e$. This follows from the fact that the top horizontal functor sends $A^{G/e}$ to a terminal object, see Lemma 4.13.
Uniqueness follows from the fact that the nullhomotopy is unique: the component of 00 in $\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Perf}(A)^{tG})$ is contractible because 00 is terminal. Alternatively, one can consider this as a square of semirings in spaces, where uniqueness is because projecting off a component amounts to inverting an idempotent.
Fully faithfulness of $\iota$ follows from Theorem 6.3 and the fact that fully faithfulness is preserved by the groupoid functor. ∎
@@ 8. Pullback decomposition of equivariant spectra
for the $\infty$-category of $R$-modules in $\operatorname{Sp}_{G}$ and $\operatorname{Sp}_{G}/\mathscr{F}$ respectively. More concretely, we have $\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F}\simeq\operatorname{Mod}_{\operatorname{Sp}_{G}}(R\otimes\widetilde{E}\mathscr{F})$. We denote by
@@
$\Phi_{R}^{K}\colon\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)\xrightarrow{\widetilde{\Phi}_{R}^{K}}\operatorname{Mod}(\Phi^{K}(R))^{hW_{G}(K)})\xrightarrow{\mathrm{fgt}}\operatorname{Mod}(\Phi^{K}(R)).$
@@
[…]
Recall the functors $\pi$ and $\beta$ from 4.4 and Lemma 4.9.
@@
${(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F})^{\omega}}$${(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F}\cup K)^{\omega}}$${\operatorname{Fun}(BW_{G}(K),\operatorname{Perf}(\Phi^{K}(R)))}$${\prod_{0<p\mid|W_{G}(K)|}\mathrm{StMod}^{\mathrm{fin}}((\Phi^{K}R)_{p}^{\wedge}W_{G}(K)).}$$\scriptstyle{L_{\mathscr{F}\cup K}^{\omega}}$$\scriptstyle{(\widetilde{\Phi}_{R}^{K})^{\omega}}$$\scriptstyle{\beta\circ(\overline{\Phi}^{K}_{R})^{\omega}}$$\scriptstyle{\beta\circ\pi}$
@@
By [NPR24, Corollary 7.13], the following is a pullback square
${(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F})^{\omega}}$${(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F}\cup K)^{\omega}}$${\operatorname{Fun}(BW_{G}(K),\operatorname{Perf}(\Phi^{K}(R)))}$${\mathrm{StMod}^{\mathrm{fin}}((\Phi^{K}R)W_{G}(K)).}$$\scriptstyle{L_{\mathscr{F}\cup K}^{\omega}}$$\scriptstyle{(\widetilde{\Phi}_{R}^{K})^{\omega}}$$\scriptstyle{(\overline{\Phi}_{R}^{K})^{\omega}}$$\scriptstyle{\pi}$
We observe that the above square differs from the one in the theorem by the bottom right corner. However, by Lemma 4.9, the functor $\beta$ relating the two vertices is fully faithful, and so the pullbacks of the two squares are equivalent. ∎
@@ 9. The classification result
which is symmetric monoidal. We let $R^{X}$ denote the image of $\mathbb{D}(\Sigma_{+}^{\infty}X)$ under this functor, and observe that $R^{X}$ is a separable commutative algebra which is compact as an $R$-module. We therefore obtain a functor of groupoids
(9.1) $R^{(-)}\colon\mathcal{F}\mathrm{in}^{\operatorname{op}}_{G}\to\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)^{\omega}).$
More generally, given a family $\mathscr{F}$ of subgroups of $G$, we have a functor
(9.2) $R^{(-)}_{/\mathscr{F}}\colon\mathcal{F}\mathrm{in}_{G}^{\operatorname{op}}/\mathscr{F}\subseteq\mathcal{F}\mathrm{in}_{G}^{\operatorname{op}}\stackrel{{\scriptstyle R^{(-)}}}{{\longrightarrow}}\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)^{\omega})\longrightarrow\mathcal{C}\mathcal{S}\mathrm{ep}((\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F})^{\omega}),$
@@
Note that Equation 9.2 generalizes Equation 9.1 because $R^{(-)}_{/\varnothing}=R^{(-)}$.
Terminology 9.3.
@@
Now assume that $K\subseteq G$ is a subgroup such that $K\not\in\mathscr{F}$ and $K^{\prime}\in\mathscr{F}$ for all $K^{\prime}\subsetneq K$. We note that the canonical inclusion $i_{K}\colon\mathcal{F}\mathrm{in}_{G}/\mathscr{F}\cup K\to\mathcal{F}\mathrm{in}_{G}/\mathscr{F}$ admits a retraction
@@
which, using the decomposition given in 7.8, sends the orbit $G/K$ to the empty set. The next result illustrates how the functors $R^{(-)}_{/\mathscr{F}}$ and $R^{(-)}_{/\mathscr{F}\cup K}$ relate to one another.
@@
[…]
@@
As already discussed in the introduction, our goal in this paper is to find conditions to ensure that the functor (Equation 9.2) is an equivalence. The next result enables an induction on $\mathcal{F}$.
@@
[…]
The right most face of the cube is obtained by applying $\mathcal{C}\mathcal{S}\mathrm{ep}(-)$ to the pullback square of symmetric monoidal stable $\infty$-categories of Theorem 8.2. In particular, this is a pullback square. Consider the left most face of the cube: this commutes because for $X\in\mathcal{F}\mathrm{in}_{G}$, the $K$-fixed points $X^{K}$ is a free $\Gamma$-set if $X\simeq\coprod_{i}G/K$ in $\mathcal{F}\mathrm{in}_{G}$. We claim that this left face of the cube is a pullback. This follows from the pasting lemma, [Lur09, Proposition 4.4.2.1], since the bottom left square and the outer square are pullback squares, for the latter use 7.8.
@@
$\simeq\iota\operatorname{proj}(i_{K}(X)^{K})$
$\simeq\iota(\pi_{K}i_{K}(X))^{K}\simeq\iota(X)^{K}$
where in the last step we used fully faithfulness of $i_{K}$.
@@
with each $\mathscr{F}_{i}\backslash\mathscr{F}_{i-1}=\{(K_{i})\}$ consisting of a single conjugacy class. Note that this implies that for all proper subgroups $K\subsetneq K_{i}$ we have $K\in\mathscr{F}_{i-1}$. We will prove by descending induction on $i\leq n-1$, that the functor
@@
is an equivalence. The case $i=0$ will then be our claim. If $i=n-1$, then $\mathscr{F}_{i}=\mathscr{F}_{n-1}$ is the family of proper subgroups, so taking $G$-geometric fixed points induces an equivalence $\Phi^{G}\colon(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F}_{n-1})^{\omega}\xrightarrow{\sim}\operatorname{Perf}(\Phi^{G}(R))$. Since $\Phi^{G}(R)$ is separably closed, we also deduce that
@@
is an equivalence. By the $2$-out-of-$3$ property we conclude that $R_{/\mathscr{F}_{n-1}}^{(-)}$ is an equivalence proving the base case. The inductive step then follows from Theorem 9.5 and so we are done by induction. ∎
@@
Let $p$ be a prime number and $G$ a $p$-group. Then there are equivalences of $\infty$-groupoids
$\mathcal{F}\mathrm{in}_{G}^{\operatorname{op}}\xrightarrow{\sim}\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Sp}_{G}^{\omega})\quad\mathrm{and}\quad\mathcal{F}\mathrm{in}_{G}^{\operatorname{op}}\xrightarrow{\sim}\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Mod}_{\operatorname{Sp}_{G}}(\mathbb{Z}_{G})^{\omega}).$
@@
If $G$ is trivial, the two claims follows from [NP24, Corollary 10.6]. If $G$ is a nontrivial $p$-group, then for all $K\subseteq G$, the $K$-geometric fixed points are $\Phi^{K}\mathbb{S}_{G}=\mathbb{S}$ and $\Phi^{K}\mathbb{Z}_{G}=\mathbb{Z}$ both with trivial Weyl group action. Note that these commutative algebras are all separably closed by [NP24, Corollary 10.6] again. They also satisfy IC and RC by Theorem 5.5, Lemma 6.5 and 5.4. So the claims follow from 9.6 applied to $\mathscr{F}=\varnothing$. ∎
@@
Let $p$ be a prime number and $G$ a $p$-group. Then the Galois groups of $\operatorname{Sp}_{G}^{\omega}$ and $\operatorname{Mod}_{\operatorname{Sp}_{G}}(\mathbb{Z}_{G})^{\omega}$ are trivial.
@@
To treat both cases simultaneously, let us write $A^{G}_{H}$ both for the separable commutative algebra $\mathbb{D}(G/H_{+})$ in $\operatorname{Sp}_{G}^{\omega}$ and for $\mathbb{Z}_{G}\otimes\mathbb{D}(G/H_{+})\in\operatorname{Mod}_{\operatorname{Sp}_{G}}(\mathbb{Z}_{G})^{\omega}$. First of all note that $\pi_{0}(\mathbb{1})$ in both categories is given by the Burnside ring $A(G)$, see [PSW22, Remark 2.20] for the second category. It follows that the two categories are indecomposable, see 10.1, and so their Balmer spectra are connected by [NP24, Proposition 3.7]. Since the finite Galois covers are exactly thoses separable algebras with locally constant (i.e. constant) degree function ([NP24, Corollary 8.13]), using 9.7 it will suffices to check that the degree function of $A^{G}_{H}$ is constant if and only if $G=H$. But in both cases the degree function is nonzero exactly at points in $\operatorname{supp}(A^{G}_{H})$ which is a proper subset of the Balmer spectrum unless $G=H$, see [PSW22, Lemma 2.12 and Proposition 2.13] and [BS17, Corollary 4.13]. ∎

### v2

@@ 7. Separably closed commutative algebras
Next, we record a consequence of this condition.
@@
[…]
@@
Construction 7.10.
For any two families $\mathscr{F}\subseteq\mathscr{F}^{\prime}$, the fully faithful inclusion $\iota\colon\operatorname{Fin}_{G}/\mathscr{F}^{\prime}\subseteq\operatorname{Fin}_{G}/\mathscr{F}$ admits a right adjoint (hence a colocalization)
$\pi\colon\operatorname{Fin}_{G}/\mathscr{F}\longrightarrow\operatorname{Fin}_{G}/\mathscr{F}^{\prime}$
with counit
$\pi(X)=\{x\in X\,|\,\mathrm{Stab}_{G}(x)\not\in\mathscr{F}^{\prime}\}\subseteq X.$
We also write $X/\mathcal{F}$ for $\pi(X)$.
Proposition 7.11.
Let $G$ be a finite group, $\mathscr{F}$ a family of subgroups of $G$ and $K\subseteq G$ a subgroup such that $K\not\in\mathscr{F}$, but for all proper subgroups $K^{\prime}\subsetneq K$ we have $K^{\prime}\in\mathscr{F}$. Then in the solid diagram
$\operatorname{Fin}_{G}/\mathscr{F}$$\operatorname{Fin}_{G}/(\mathscr{F}\cup K)$$\operatorname{Fin}_{W_{G}(K)}$$\operatorname{Fin}_{W_{G}(K)}/\{e\}$$\scriptstyle{\pi}$$\scriptstyle{(-)^{K}}$$\scriptstyle{\Phi}$$\scriptstyle{\pi^{\prime}}$
there is a unique factorization $\Phi$ of $\pi^{\prime}\circ(-)^{K}$ through $\pi$. In fact $\Phi$ is given by the composite
$\operatorname{Fin}_{G}/(\mathscr{F}\cup K)\xrightarrow{\iota}\operatorname{Fin}_{G}/\mathscr{F}\xrightarrow{(-)^{K}}\operatorname{Fin}_{W_{G}(K)}\xrightarrow{\pi^{\prime}}\operatorname{Fin}_{W_{G}(K)}/\{e\}.$
We note that while both horizontal functors have left adjoints, this square is generally not left adjointable. In fact, it is exactly when it is a pullback square, which happens when every subgroup has a strictly larger normalizer in $G$. It however follows from 7.8 that it is always a pullback when restricting to maximal subgroupoids.
Proof.
Since $\pi$ is a colocalization, the claim about the factorization is equivalent to the functor $\pi^{\prime}\circ(-)^{K}$ inverting the counit
$\iota\colon\pi(X)=\{x\in X\,|\,\mathrm{Stab}_{G}(x)\not\in\mathscr{F}\cup K\}\subseteq X,\,X\in\operatorname{Fin}_{G}/\mathscr{F}.$
Now, $\pi^{\prime}(\iota^{K})$ computes to be the inclusion
$\{x\in X^{K}\,|\,\mathrm{Stab}_{G}(x)\not\in\mathscr{F}\cup K\mbox{ and }\mathrm{Stab}_{W_{G}(K)}(x)\neq\{e\}\}\subseteq\{x\in X^{K}\,|\,\mathrm{Stab}_{W_{G}(K)}(x)\neq\{e\}\},$
which is immediately checked to be an equality. The resulting formula for $\Phi$ is a formal consequence of the above discussion. ∎
@@
Let $G$ be a finite group and assume $A\in\operatorname{Fun}(BG,\operatorname{CAlg})$ satisfies IC and RC. Then, there is a unique factorization $\iota$
$\operatorname{CAlg}^{\operatorname{sep}}(\operatorname{Perf}(A)^{hG})$$\operatorname{CAlg}^{\operatorname{sep}}(\operatorname{Perf}(A)^{tG})$$\operatorname{Fin}^{\operatorname{op}}_{G}$$\left(\operatorname{Fin}_{G}/\{e\}\right)^{\operatorname{op}}$$\scriptstyle{q}$$\scriptstyle{A^{(-)}}$$\scriptstyle{\pi}$$\scriptstyle{\mathrm{\iota}}$
@@
To get the factorization, we need to see that $q\circ A^{(-)}$ inverts the counit
$\pi(X)=\{x\in X\,|\,\mathrm{Stab}_{G}(x)\neq\{e\}\}\subseteq X\,,\,X\in\operatorname{Fin}_{G}.$
This is clear since $q(A^{G})=0$. Full faithfulness of $\iota$ follows from Theorem 6.3. ∎
@@ 8. Pullback decomposition of equivariant spectra
[…]
@@
$\Phi_{R}^{K}\colon\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)\xrightarrow{\widetilde{\Phi}_{R}^{K}}\operatorname{Mod}(\Phi^{K}(R))^{hW_{G}(K)}\xrightarrow{\mathrm{fgt}}\operatorname{Mod}(\Phi^{K}(R)).$
@@
$(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathcal{F})^{\omega}$$(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathcal{F}\cup K)^{\omega}$$\operatorname{Perf}(\Phi^{K}R)^{hW_{G}(K)}$$\operatorname{Perf}(\Phi^{K}R)^{tW_{G}(K)}.$$\scriptstyle{(\widetilde{\Phi}_{R}^{K})^{\omega}}$$\scriptstyle{L^{\omega}_{\mathcal{F}\cup K}}$$\scriptstyle{(\overline{\Phi}^{K}_{R})^{\omega}}$$\scriptstyle{q}$
@@
$(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F})^{\omega}$$(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F}\cup K)^{\omega}$$\operatorname{Fun}(BW_{G}(K),\operatorname{Perf}(\Phi^{K}(R)))$$\prod_{0<p\mid|W_{G}(K)|}\operatorname{Perf}(\Phi^{K}(R)^{\wedge}_{p})^{tW_{G}(K)}$$\scriptstyle{L_{\mathscr{F}\cup K}^{\omega}}$$\scriptstyle{(\widetilde{\Phi}_{R}^{K})^{\omega}}$$\scriptstyle{\beta_{0}\circ(\overline{\Phi}^{K}_{R})^{\omega}}$$\scriptstyle{\beta_{0}\circ q}$
where $\beta_{0}$ is induced by the base-change functors.
@@
We claim that the base-change functors assemble into a fully faithful functor
$\beta_{0}\colon\operatorname{Perf}(\Phi^{K}R)^{tW_{G}(K)}\to\prod_{p}\operatorname{Perf}((\Phi^{K}R)^{\wedge}_{p})^{tW_{G}(K)}.$
Indeed the above functor can be obtained from Lemma 4.9(b) by applying the idempotent completion functor, and by identifying the result with the Tate categories via 4.14. We observe that the square in the corollary differs from Theorem 8.2 by the bottom right corner. Since we have a fully faithful functor relating the two different vertices, we obtain the pullback of the two squares are equivalent, proving the corollary. ∎
@@ 9. The classification result
which is symmetric monoidal. We let $R^{X}$ denote the image of $\mathbb{D}(\Sigma_{+}^{\infty}X)$ under this functor, and observe that $R^{X}$ is a separable commutative algebra which is compact as an $R$-module. We therefore obtain a functor
(9.1) $R^{(-)}\colon\operatorname{Fin}^{\operatorname{op}}_{G}\to\operatorname{CAlg}^{\operatorname{sep}}(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)^{\omega}).$
More generally, given a family $\mathscr{F}$ of subgroups of $G$, we have a functor $R^{(-)}_{/\mathscr{F}}$:
(9.2) $(\operatorname{Fin}_{G}/\mathscr{F})^{\operatorname{op}}\subseteq\operatorname{Fin}_{G}^{\operatorname{op}}\stackrel{{\scriptstyle R^{(-)}}}{{\longrightarrow}}\operatorname{CAlg}^{\operatorname{sep}}(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)^{\omega})\longrightarrow\operatorname{CAlg}^{\operatorname{sep}}((\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F})^{\omega}),$
@@
Note that (9.2) generalizes (9.1) because $R^{(-)}_{/\varnothing}=R^{(-)}$. We now give conditions under which this functor is fully faithful.
Proposition 9.3.
Let $G$ be a finite group and $\mathscr{F}$ a family of subgroups of $G$. Consider $R\in\operatorname{CAlg}(\operatorname{Sp}_{G})$ and suppose that for all subgroups $H\subseteq G$ with $H\not\in\mathscr{F}$, the commutative ring $\pi_{0}^{H}(R)$ is indecomposable, and $\pi_{0}(\Phi^{H}R)$ is nonzero. Then the functor (9.2) is fully faithful.
Proof.
We argue essentially as in the proof of Theorem 6.3, namely we work by induction on $G$. We consider the map
$\operatorname{Map}_{\operatorname{Fin}_{G}/\mathscr{F}}(Y,X)\to\operatorname{Map}_{\operatorname{CAlg}(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F})}(R^{X},R^{Y})$
which we want to show is an equivalence. Since our functor sends coproducts to products, we may assume without loss of generality that $Y$ is an orbit, say $G/H$, with $H\notin\mathscr{F}$. We then have compatible equivalences
$\operatorname{Map}_{\operatorname{Fin}_{G}/\mathscr{F}}(Y,X)\simeq\operatorname{Map}_{\operatorname{Fin}_{H}/\mathscr{F}_{H}}(*,\operatorname{res}^{G}_{H}(X))$
and
$\operatorname{Map}_{\operatorname{CAlg}(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R))/\mathscr{F})}(R^{X},R^{Y})\simeq\operatorname{Map}_{\operatorname{CAlg}(\operatorname{Mod}_{\operatorname{Sp}_{H}}(R)/\mathscr{F}_{H})}(R^{\operatorname{res}^{G}_{H}(X)},R)$
where $\mathscr{F}_{H}$ is the family of subgroups of $H$ in $\mathscr{F}$. We may assume that $\operatorname{res}^{G}_{H}(X)$ has no isotropy in $\mathscr{F}_{H}$ since if $X_{0}\subset X$ is the subset of points with isotropy not in $\mathscr{F}_{H}$, the map $\operatorname{Map}_{\operatorname{Fin}_{G}}(Y,X_{0})\to\operatorname{Map}_{\operatorname{Fin}_{G}}(Y,X)$ is a bijection (any point in the image of a map $Y\to X$ has isotropy conjugate to a group containing $H$, and $H$ is not in $\mathscr{F}$). Thus, by induction we are reduced to the case Y=Y=*. Now, since $\pi_{0}^{G}(R)$ is indecomposable, we may also assume that $X$ is an orbit (since coproducts of $X$’s pull out on either side to coproducts). Now if $X=G/G$, both sides are contractible since $R$ is the unit in $\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)$, and if $X=G/H$ for $H$ a proper subgroup, both sides are empty, since $\Phi^{G}(R^{X})=0$ and $\Phi^{G}(R)$ is nonzero. ∎
In order to discuss when the functor (9.2) is essentially surjective, it will be useful to consider the induced functor on groupoids
(9.4) $R^{(-)}_{/\mathscr{F}}\colon(\mathcal{F}\mathrm{in}_{G}/\mathscr{F})^{\operatorname{op}}\longrightarrow\mathcal{C}\mathcal{S}\mathrm{ep}((\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F})^{\omega}),$
Terminology 9.5.
@@
Now assume that $K\subseteq G$ is a subgroup such that $K\not\in\mathscr{F}$ and $K^{\prime}\in\mathscr{F}$ for all $K^{\prime}\subsetneq K$. Recall from 7.10 that the canonical inclusion $i_{K}\colon\mathcal{F}\mathrm{in}_{G}/\mathscr{F}\cup K\to\mathcal{F}\mathrm{in}_{G}/\mathscr{F}$ admits a right adjoint at the level of categories, and hence we get a retraction at the level of groupoids:
@@
Definitionally, it sends the orbit $G/K$ to the empty set.
@@
$\mathcal{F}\mathrm{in}_{G}^{\operatorname{op}}/\mathscr{F}$$\mathcal{F}\mathrm{in}_{G}^{\operatorname{op}}/\mathscr{F}\cup K$$\mathcal{C}\mathcal{S}\mathrm{ep}((\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F})^{\omega})$$\mathcal{C}\mathcal{S}\mathrm{ep}((\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F}\cup K)^{\omega}).$$\scriptstyle{R_{/\mathscr{F}}^{(-)}}$$\scriptstyle{\pi_{K}^{\operatorname{op}}}$$\scriptstyle{R^{(-)}_{/\mathscr{F}\cup K}}$$\scriptstyle{L_{\mathscr{F}\cup K}}$
@@
As already discussed in the introduction, our goal in this paper is to find conditions to ensure that the functor (9.4) is an equivalence. The next result enables an induction on $\mathcal{F}$.
@@
$(\mathcal{F}\mathrm{in}_{G}/\mathscr{F}\cup K)^{\mathrm{op}}$$\mathcal{C}\mathcal{S}\mathrm{ep}((\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathcal{F}\cup K)^{\omega})$$(\mathcal{F}\mathrm{in}_{G}/\mathscr{F})^{\mathrm{op}}$$\mathcal{C}\mathcal{S}\mathrm{ep}((\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathcal{F})^{\omega})$$\prod_{e\not=(U)\subseteq\Gamma}(\mathcal{F}\mathrm{in}_{W_{\Gamma}(U)}^{\mathrm{free}})^{\mathrm{op}}$$\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Perf}(\Phi^{K}R)^{t\Gamma})$$\mathcal{F}\mathrm{in}_{\Gamma}^{\mathrm{op}}\simeq{\prod_{(U)\subseteq\Gamma}(\mathcal{F}\mathrm{in}_{W_{\Gamma}(U)}^{\mathrm{free}})^{\mathrm{op}}}$$\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Perf}(\Phi^{K}R)^{h\Gamma})$$\ast$$(\mathcal{F}\mathrm{in}_{\Gamma}^{\mathrm{free}})^{\mathrm{op}}$$\scriptstyle{R_{/\mathscr{F}\cup K}^{(-)}}\scriptstyle{\sim}$$\scriptstyle{\overline{(-)^{K}}}$$\scriptstyle{\overline{\Phi}^{K}}$$\scriptstyle{\pi_{K}}$$\scriptstyle{R_{/\mathscr{F}}^{(-)}}$$\scriptstyle{(-)^{K}}$$\scriptstyle{L_{\mathscr{F}\cup K}}$$\scriptstyle{\Phi^{K}}$$\scriptstyle{\iota}$$\scriptstyle{\mathrm{proj}}$$\scriptstyle{(\Phi^{K}R)^{(-)}}\scriptstyle{\sim}$$\scriptstyle{\mathrm{proj}}$$\scriptstyle{q}$
The right most face of the cube is obtained by applying $\mathcal{C}\mathcal{S}\mathrm{ep}(-)$ to the pullback square of symmetric monoidal stable $\infty$-categories of Theorem 8.2. In particular, this is a pullback square. The left face of the cube is obtain from Proposition 7.11 by passing to groupoids. In particular, the square is commutative and the functor $\overline{(-)^{K}}$ is given by the composite $\operatorname{proj}((i_{K}(-))^{K})$. We now claim that this left face of the cube is a pullback. This follows from the pasting lemma, [LUR09, Proposition 4.4.2.1], since the bottom left square and the outer left square are pullback squares, for the latter use 7.8.
@@
$\simeq\iota\operatorname{proj}(i_{K}(X)^{K})=\iota(\overline{X^{K}})$
where in the last step we used the definition of $\overline{(-)^{K}}$.
@@
[…]
@@
is an equivalence. The case $i=0$ will then be our claim. If i=n1i=n-1, then $\mathscr{F}_{i}=\mathscr{F}_{n-1}$ is the family of proper subgroups, so taking $G$-geometric fixed points induces an equivalence $\Phi^{G}\colon(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F}_{n-1})^{\omega}\xrightarrow{\sim}\operatorname{Perf}(\Phi^{G}(R))$. Since $\Phi^{G}(R)$ is separably closed, we also deduce that
@@
is an equivalence. By the 22-out-of-33 property we conclude that $R_{/\mathscr{F}_{n-1}}^{(-)}$ is an equivalence proving the base case. The inductive step then follows from Theorem 9.7 and so we are done by induction. ∎
@@
Let $p$ be a prime number and $G$ a $p$-group. Then the functor (9.1) induce equivalences
$\operatorname{Fin}_{G}^{\operatorname{op}}\xrightarrow{\sim}\operatorname{CAlg}^{\operatorname{sep}}(\operatorname{Sp}_{G}^{\omega})\quad\mathrm{and}\quad\operatorname{Fin}_{G}^{\operatorname{op}}\xrightarrow{\sim}\operatorname{CAlg}^{\operatorname{sep}}(\operatorname{Mod}_{\operatorname{Sp}_{G}}(\mathbb{Z}_{G})^{\omega}).$
@@
If $G$ is a $p$-group, then for all $K\subseteq G$, the $K$-geometric fixed points are $\Phi^{K}\mathbb{S}_{G}=\mathbb{S}$ and $\Phi^{K}\mathbb{Z}_{G}=\mathbb{Z}$ both with trivial Weyl group action. Note that these commutative algebras are all separably closed by [NP24, Corollary 10.6] again. They satisfy IC by Theorem 5.5 and 5.4, and RC by 6.6. Thus the fact that the above functors are essentially surjective follow from 9.8 applied to $\mathscr{F}=\varnothing$. For fully faithfulness, we apply Proposition 9.3 and note that for any subgroups $H\subseteq G$, the groups $\pi_{0}^{H}(\operatorname{\mathbb{S}}_{H})$ and $\pi_{0}^{H}(\mathbb{Z}_{H})$ agree with the Burnside ring $A(H)$, see [PSW22, Remark 2.20] for the second category, which is indecomposable by a result of Dress, see 10.1. ∎
@@
Let $p$ be a prime number and $G$ a $p$-group. Then the Galois groups of $\operatorname{Sp}_{G}$ and $\operatorname{Mod}_{\operatorname{Sp}_{G}}(\mathbb{Z}_{G})$ are trivial.
@@
To treat both cases simultaneously, let us write $A^{G}_{H}$ both for the separable commutative algebra $\mathbb{D}(G/H_{+})$ in $\operatorname{Sp}_{G}^{\omega}$ and for $\mathbb{Z}_{G}\otimes\mathbb{D}(G/H_{+})\in\operatorname{Mod}_{\operatorname{Sp}_{G}}(\mathbb{Z}_{G})^{\omega}$. We have already noted that $\pi_{0}(\mathbb{1})$ in both categories is given by the Burnside ring $A(G)$ which is indecomposable. It follows that their Balmer spectra are connected by [NP24, Proposition 3.7]. Since the finite Galois covers are exactly those separable algebras with underlying compact object and with locally constant (i.e. constant) degree function ([NP24, Corollary 8.13]), using 9.9 it will suffice to check that the degree function of $A^{G}_{H}$ is constant if and only if $G=H$. But in both cases the degree function is nonzero exactly at points in $\operatorname{supp}(A^{G}_{H})$ which is a proper subset of the Balmer spectrum unless $G=H$, see [PSW22, Lemma 2.12 and Proposition 2.13] and [BS17, Corollary 4.13]. ∎